13 24 1of the Normal table. Click here to visw 1. What is the P-value? The P value is (Round to four decimal places as needed ) g. What are the critical value(s)? (Round to three decimal places as needed. Use a comma to separate answers as needed) h. What is the area of the critical region? Click to select your answeris). 6/ 0
Answer #1

Claim : To test whether that when the parents used the particular method of gender selection then the proportion of baby girls is different from 0.5 or not

Hypothesis : o : p = 0.5

H1 : p \neq 0.5

Two tailed test

Given that

sample size = n = 144

Let x = Total number of girls in the sample = 78

Therefore, sample proportion = \hat p = \frac{x}{n} = \frac{78}{144} = \frac{13}{24}=0.5417

Test statistics : z=\frac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}}

z=\frac{0.5417-0.5}{\sqrt{\frac{0.5*(1-0.5)}{144}}}=1

z=1

Pvalue Can be calculated as

Pvalue = 2 * ( 1 - P(|Z| < z ) )

Pvalue = 2 * ( 1 - P(|Z| < 1 ) )

Pvalue = 2 * ( 1 - 0.8413 ) ...............( From Z table )

= 2 * 0.1586 = 0.3173

Pvalue = 0.3173 .............(ANSWER )

====================================================================

G ) Critical value : Zc = -1.64 OR Zc = 1.64 ............( From Z table )

=========================================================================

H ) Area of the critical region is

-Zc--1.64 2c 1.64